OpenStudy (anonymous):

Find csc x if sin x + cot x cos x = square root 3

OpenStudy (anonymous):

express cot(x) as cos(x)/sin(X) and simplify :)

OpenStudy (anonymous):

i'm still confused, i got cos^2 x = square root 3

OpenStudy (anonymous):

you should get sin(x) as the common denominator

OpenStudy (mertsj):

\[\sin x+\frac{\cos x}{\sin x}\times \cos x=\sqrt{3}\]

OpenStudy (anonymous):

\[\sin x + \cos ^{2} x/\sin x = \sqrt[2]{3}\]

OpenStudy (anonymous):

that ^

OpenStudy (mertsj):

\[\frac{\sin ^2x+\cos ^2x}{\sin x}=\sqrt{3}\]

OpenStudy (mertsj):

\[\frac{1}{\sin x}=\sqrt{3}\]

OpenStudy (anonymous):

oohh!! thanks! I got it!

OpenStudy (mertsj):

Take reciprocals of both sides and solve